### v1

@@ Abstract
The standard mixture modelling framework has been widely used to study heterogeneous populations, by modelling them as being composed of a finite number of homogeneous sub-populations. However, the standard mixture model assumes that each data point belongs to one and only one mixture component, or cluster, but when data points have fractional membership in multiple clusters this assumption is unrealistic. It is in fact conceptually very different to represent an observation as partly belonging to multiple groups instead of belonging to one group with uncertainty. For this purpose, various soft clustering approaches, or individual-level mixture models, have been developed. In this context, Heller et al (2008) formulated the Bayesian partial membership model (PM) as an alternative structure for individual-level mixtures, which also captures partial membership in the form of attribute-specific mixtures, but does not assume a factorization over attributes.
@@ 1 Introduction
Model-based clustering has been widely used among researchers to study heterogeneous populations, by modelling them as being composed of a finite number of homogeneous sub-populations (eg. Fraley and Raftery, 2002; McLachlan and Peel, 2000; Bouveyron et al, 2019; Gormley et al, 2023). In this framework, observations within a dataset are modeled as originating from one of multiple probability distributions. The objective is to achieve a clustering solution where observations are partitioned into distinct groups. Notably, observations that show a significant posterior probability of belonging to more than one cluster are considered to have uncertain group membership. This uncertainty can sometimes suggest a model that does not adequately fit the data. However, the standard mixture model assumes that each data point belongs to one and only one mixture component, or cluster, but when data points have fractional membership in multiple clusters this assumption is unrealistic.
In contrast, mixed membership and partial membership models accommodate partial membership to multiple clusters. For example, let’s consider a football player with the role of a trequartista. His duties on the pitch change depending on the phase of play. In an offensive phase, they are those associated with the attacking role, while in a defensive phase they are those associated with the midfielder role. That football player should be represented as partly belonging to two different classes or sets. Being certain that a player’s role is partly midfielder and partly striker is very different from being uncertain about a player’s role on the pitch.
@@
Partial membership models (Heller et al, 2008) are defined using a similar framework, but they overcome some of the drawbacks of mixed membership models. In the present paper we specify partial membership models for count data, in particular when component distributions are Poisson distributions.
Eventhough partial membership models exist for some time, they are still under utilised in literature. Gruhl and Erosheva (2013) apply partial membership model to NBA player data from the 2010–11 season and compare the results with those achieved with the mixed membership model. Chen et al (2017) used partial membership for semantic image segmentation, while Hou-Liu and Browne (2022) start from partial membership models and epistatic clustering (Zhang, 2013) to develop an “hybrid method” between the two, to cluster hybrid species of flowers, which tend to exhibit a mixture of parent characteristics.
We apply the PM model to Serie A football player performance data, from the 2022/2023 football season. The aim is to estimate the various roles players are inclined to occupy on the field, beyond their primary positions, by analyzing their playing styles. The goal of the application within sports analysis is to assist coaches, talent scouts, team managers, and analysts in making more informed decisions. These decisions pertain to team strategy, talent acquisition, and advancing statistical research in sports.
@@
The article is organised as follows. Section 2 outlines the Bayesian partial membership model specification with Poisson component distributions, and compares the data generative process to those in mixed membership and mixture models. It also gives an overview on technical aspects like label switching and model selection, justifying the choice of the information criterion through literature and a simulation. In Section 3, both partial and mixed membership models are applied to Serie A football players performance data and the results are compared. For sake of completeness, a Poisson mixture model is also applied to this study. However, the results are only briefly explored, because a crisp clustering approach does not align with the specific objectives and purposes of the proposed application. Conclusions and future developments follow in Section 4. To further demonstrate the versatility and effectiveness of the proposed model, we include an additional application of the model to Washington DC bike sharing data in Appendix A.
2 Partial Membership Model
@@ 2.1 Mixture Model
where $\bm{\pi}=\{{\pi}_{k}:k=1,2,\dots,K\}$ are the mixing proportions and $\bm{\Theta}=\{\bm{\theta}_{k}:k=1,2,\ldots,K\}$ are the mixture component parameters. In this framework, a data point is assumed to come from one (and only one) of the $K$ mixture components. Let $\bm{\tau}_{i}$ be the latent component membership indicator variable for observation $i$, so $\tau_{ik}\in\{0,1\}$ and $\sum_{k}\tau_{ik}=1$. In the mixture model, we assume that $\bm{\tau}_{i}\sim\mbox{multinomial}(1,\bm{\pi})$, and so $\bm{\tau}_{i}=(\tau_{i1},\dots,\tau_{iK})$ is the binary membership vector, where $\tau_{ik}=1$ indicates that the data point $\bm{x}_{i}$ was generated by mixture component $k$.
@@ 2.2 Bayesian Partial Membership Model
In a partial membership model, we relax the constraint $\tau_{ik}\in\{0,1\}$ to take any continuous value in the range $[0,1]$ and where $\sum_{k}\tau_{ik}=1$; therefore, $\tau_{ik}$ is in the unit simplex, allowing multiple cluster memberships for each data point. In the partial membership model, the density of an observation becomes,
@@
We integrate over all values of $\bm{\tau}_{i}$ instead of summing. The product over $k$ reflects that the contribution to the likelihood of $\bm{x}_{i}$ from each component is compounded multiplicatively, representing how each component contributes to explaining $\bm{x}_{i}$ given its partial membership weights $\tau_{ik}$. The exponent $\tau_{ik}$ on $P_{k}(\bm{x}_{i}\mid\bm{\theta}_{k})$ represents the degree of membership of $\bm{x}_{i}$ in component $k$. This is not directly present in the finite mixture model, where the component membership is implicit in the mixing weights $\pi_{k}$ and assumed to be binary (fully in one component). In the partial membership context, $\tau_{ik}$ as an exponent softens this assumption, allowing $\bm{x}_{i}$ to “partially belong” to multiple components simultaneously, each to varying degrees expressed by $\tau_{ik}$.
@@ 2.3 Mixed Membership Model
A method for fitting mixed membership models to count data is outlined in White and Murphy (2016). Broadly speaking, mixed membership (MM) models operate under the assumption that each data feature (such as pixels in an image analysis scenario) of a given item (e.g., an image) is independently drawn from a mixture distribution defined by the membership vector of the item, $x_{ij}\mid\bm{\tau}_{i}\sim\sum_{k}\tau_{ik}P({x}_{ij}\mid\lambda_{kj})$ for $j=1,2,\ldots,J$. MM models make sense when the items being analyzed, such as images, can be seen as collections of exchangeable parts (e.g., pixels). In contrast, partial membership models do not rely on the assumption of exchangeable sub-objects.
@@ 2.5 Identifiability
Parameters in a model are not identified if the same likelihood function is obtained for more than one choice of the model parameters (Teicher, 1963; Gelman et al, 2013).
@@
The issue of identifiability in finite mixture distributions demands careful consideration, particularly when conducting Bayesian inference. Frühwirth-Schnatter (2006) describes three types of non-identifiability: invariance to relabeling of the components in the mixture model, non-identifiability due to potential overfitting, and a generic non-identifiability in certain classes of mixture distributions. Non identifiability can lead to poor convergence of MCMC methods used in Bayesian inference.
To draw valid posterior inference, the invariance to relabeling characteristic requires particular attention. Among the possible solutions, such as Celeux et al (2000) or Murphy et al (2020), in this paper the label switching problem has been addressed permuting after the run of the model the labels to each MCMC draw, using the probabilistic relabelling algorithm of Sperrin et al (2010), provided by the R package $\mathsf{label.switching}$.
In the application proposed in the present paper, potential identifiability challenges are addressed through the incorporation of entities termed as archetypal units. These units emerge from a statistical technique known as archetypal analysis. The primary objective of archetypal analysis is to identify a limited number of extreme, yet possibly hypothetical, data points (referred to as the archetypal units) within a multivariate dataset. The essence of this method is that all observed data points can be effectively represented as convex combinations of these archetypes, which form what it is called the “performance profile”. These archetypes are themselves constrained to be convex combinations of actual data points and lie on the data set boundary, i.e., the convex hull. This statistical method was first introduced by Cutler and Breiman (1994) and has found applications in different areas, e.g., in economics (Porzio et al, 2008), astrophysics (Chan et al, 2003), pattern recognition (Bauckhage and Thurau, 2009) and sports analytics (Eugster, 2012; Seth and Eugster, 2016).
In addressing the identifiability issue within the PM model, we employed archetypal analysis, guided by the presence of units who distinctly belong to a single cluster. These units, who exhibit clear archetype-defining characteristics of a particular cluster, provide anchor points. These anchors not only crystallize the definition of each cluster but also enhance the overall stability and interpretability of the model. They serve as concrete examples against which the characteristics of more ambiguously classified units can be compared, offering a benchmark for cluster assignment. Moreover, the presence of such well-defined units aids in delineating the boundaries of each cluster. This clarity is crucial in a partial membership context, where the risk of overly fluid cluster definitions could lead to diminished utility of the model. By having these distinctly categorized units, namely football players in the proposed application, we maintain a balance in our model; it remains flexible enough to accommodate the partial memberships of players who exhibit mixed attributes, yet it is anchored firmly enough to prevent the clusters from becoming indistinct or overlapping excessively.
@@
We choose a uniform distribution with parameters $(0,10)$ as the hyperprior for the $\bm{\delta}$ parameter in the mixture weights from a Dirichlet distribution: $\tau_{i}\sim\mbox{Dirichlet}(\bm{\delta})$. This decision was instrumental in ensuring the identification of archetypal players, while preserving the interpretability of the results presented in Section 3. It is important to note that for sake of completeness, we experimented with increasing the sparsity of the mixture weights matrix by opting for a smaller scalar parameter for the Dirichlet distribution (Wang and Blei, 2009). However, this increased sparsity led to a higher optimal number of clusters according to the information criterion used. Consequently, the uniform distribution emerged as our preferred choice. It effectively facilitated the identification of archetypal units while maintaining a manageable number of clusters, thus ensuring the results remained interpretable.

### v2

@@ Abstract
The standard mixture modeling framework has been widely used to study heterogeneous populations, by modeling them as being composed of a finite number of homogeneous sub-populations. However, the standard mixture model assumes that each data point belongs to one and only one mixture component, or cluster, but when data points have fractional membership in multiple clusters this assumption is unrealistic. It is in fact conceptually very different to represent an observation as partly belonging to multiple groups instead of belonging to one group with uncertainty. For this purpose, various soft clustering approaches, or individual-level mixture models, have been developed. In this context, Heller et al (2008) formulated the Bayesian partial membership model (PM) as an alternative structure for individual-level mixtures, which also captures partial membership in the form of attribute-specific mixtures.
@@ 1 Introduction
Model-based clustering has been widely used among researchers to study heterogeneous populations, by modeling them as being composed of a finite number of homogeneous sub-populations (eg. Fraley and Raftery, 2002; McLachlan and Peel, 2000; Bouveyron et al, 2019; Gormley et al, 2023). In this framework, observations within a dataset are modeled as originating from one of multiple probability distributions. The objective is to achieve a clustering solution where observations are partitioned into distinct groups. Notably, observations that show a significant posterior probability of belonging to more than one cluster are considered to have uncertain group membership. This uncertainty can sometimes suggest a model that does not adequately fit the data. However, the standard mixture model assumes that each data point belongs to one and only one mixture component or cluster, but when data points have fractional membership in multiple clusters this assumption is unrealistic. In contrast, mixed membership and partial membership models accommodate partial membership to multiple clusters. In a standard finite mixture model, if an observation appears to have a 0.5 probability of belonging to two clusters, this is interpreted as uncertainty about which single cluster is correct. By contrast, in a partial membership model, such a 50/50 split indicates that the observation is genuinely comprised of both clusters to an equal degree, rather than being purely one or the other. This conceptual difference is central to the idea of partial membership. For example, let’s consider a football player with the role of playmaker. His duties on the pitch change depending on the phase of play. In an offensive phase, they are those associated with the attacking role, while in a defensive phase they are those associated with the midfielder role. That football player should be represented as partly belonging to two different classes or sets. Being certain that a player’s role is partly midfielder and partly striker is very different from being uncertain about a player’s role on the pitch.
@@
Partial membership models (Heller et al, 2008) are defined using a similar framework, but they overcome some of the drawbacks of mixed membership models, such as the assumption of conditional independence between attributes and the challenges associated with interpreting cluster memberships in complex datasets. By relaxing these assumptions, partial membership models provide a more flexible and interpretable framework, particularly in applications with nuanced relationships between attributes, such as football player performance analysis. While Heller et al (2008) focus on continuous data and a more general setup, in the present paper, we build on the partial membership framework by adapting it specifically for count data and incorporating independent Poisson likelihoods (with Gamma priors) for each variable within a cluster. This approach is well-suited to discrete, nonnegative observations. We also address potential identifiability concerns by leveraging “archetypal units,” i.e., observations whose membership vector $\bm{\tau}_{i}$ is very close to 1 for one cluster. By identifying real examples that typify each cluster’s core characteristics, this step facilitates both interpretability and stable estimation of partial memberships.
Even though partial membership models exist for some time, they are still under utilized in literature. Gruhl and Erosheva (2013) apply partial membership model to NBA player data from the 2010–11 season and compare the results with those achieved with the mixed membership model. Chen et al (2017) used partial membership for semantic image segmentation, while Hou-Liu and Browne (2022a) starts from partial membership models and epistatic clustering (Zhang, 2013) to develop a “hybrid method” between the two, to cluster hybrid species of flowers, which tend to exhibit a mixture of parent characteristics. An alternative approach to modeling multiple cluster structures is presented by Galimberti and Soffritti (2007), who propose a method that partitions variables into disjoint subsets, applying separate mixture models to each subset. This allows an observation to belong to different clusters based on varying subsets of variables, effectively capturing multiple clustering patterns within the data. In contrast, our Partial Membership model assigns each observation a membership vector $\tau$ over the entire set of variables, providing an integrated representation of partial memberships across all clusters simultaneously. This unified approach facilitates straightforward interpretation in contexts where a holistic understanding of cluster memberships is desired. Galimberti and Soffritti (2007) method is instead useful when distinct clustering structures exist in different subsets of variables. Additionally, Hou-Liu and Browne (2022b) propose a factor and hybrid components approach for Gaussian model-based clustering, which shares a conceptual resemblance to partial membership in that some parameters can be “hybridized” across components.
To demonstrate the practical utility of the extended PM model, we apply the PM model to Serie A football player performance data, from the 2022/2023 football season. The aim is to estimate the various roles players are inclined to occupy on the field, beyond their primary positions, by analyzing their playing styles. The goal of the application within sports analysis is to assist coaches, talent scouts, team managers, and analysts in making more informed decisions. These decisions pertain to team strategy, talent acquisition, and advancing statistical research in sports. Another application presented in the Appendix A is on the daily usage of bike sharing bikes in Washington, from June 15th to July 15th, 2022. For both the bike-sharing and football scenarios, the variables of interest (number of rides originating at each station, number of shots, passes, tackles, etc.) are nonnegative integer counts with no intrinsic upper bound. The Poisson distribution is the most natural baseline model for count data under these conditions and is widely used across sports analytics and transportation research. While we acknowledge that real-world data may exhibit overdispersion or zero inflation in some contexts, the Poisson assumption provides a straightforward, interpretable starting point. Moreover, partial membership allows each observation to mix different Poisson “profiles,” thereby accommodating some variation that might otherwise lead to modest overdispersion. In cases where greater flexibility is required, extensions, such as negative binomial or zero-inflated formulations, could be explored in future work.
@@
The article is organised as follows. Section 2 outlines the modeling framework. From mixture model, to our Bayesian partial membership model specification with Poisson component distributions, and mixed membership model. It also compares the data generative process of the three models, and gives an overview on technical aspects like label switching and model selection, justifying the choice of the information criterion through literature and a simulation. In Section 3, partial membership model is applied to Serie A football players performance data and the results are compared with those achieved with the mixed membership mixture model. For the sake of completeness, a Poisson mixture model is also applied to this study, and the results are compared with those of the other two models, even if a crisp clustering approach does not align with the specific objectives and purposes of the proposed application. Conclusions and future developments follow in Section 4. To further demonstrate the versatility and effectiveness of the proposed model, we include an additional application of the model to Washington DC bike sharing data in the Appendix A.
2 Modeling Framework
@@ 2.1 Mixture Model
where the mixing proportion $\bm{\pi}=\{{\pi}_{k}:k=1,2,\dots,K\}$ satisfy $\sum_{k=1}^{K}\pi_{k}=1$ with $\pi_{k}\geq 0$. The component-specific parameters are collected in $\bm{\Theta}=\{\bm{\theta}_{k}:k=1,2,\ldots,K\}$. In this framework, each data point $\bm{x}_{i}$ is assumed to come from one (and only one) of the $K$ mixture components. Let $\bm{\tau}_{i}$ be the latent component membership indicator variable for observation $i$, so $\tau_{ik}\in\{0,1\}$ and $\sum_{k}\tau_{ik}=1$. In the mixture model, we assume that $\bm{\tau}_{i}\sim\mbox{multinomial}(1,\bm{\pi})$, and so $\bm{\tau}_{i}=(\tau_{i1},\dots,\tau_{iK})$ is the binary membership vector, where $\tau_{ik}=1$ indicates that the data point $\bm{x}_{i}$ was generated by mixture component $k$.
@@ 2.2 Bayesian Partial Membership Model
In a partial membership model, we relax the constraint $\tau_{ik}\in\{0,1\}$ to take any continuous value in the range $[0,1]$. $\tau_{i}=(\tau_{i1},\tau_{i2},\dots,\tau_{iK})$ is the vector of partial membership weights for the $i$-th observation. These weights lie in the unit simplex, such that $\tau_{ik}\geq 0$ and $\sum_{k}\tau_{ik}=1$. In the partial membership model, the density of an observation becomes,
@@
We integrate over all values of $\bm{\tau}_{i}$ instead of summing. This formulation allows for continuous membership across multiple clusters, contrasting with finite mixtures where membership is binary. The product over $k$ reflects that the contribution to the likelihood of $\bm{x}_{i}$ from each component is compounded multiplicatively, representing how each component contributes to explaining $\bm{x}_{i}$ given its partial membership weights $\tau_{ik}$. The exponent $\tau_{ik}$ on $P_{k}(\bm{x}_{i}\mid\bm{\theta}_{k})$ represents the degree of membership of $\bm{x}_{i}$ in component $k$. This is not directly present in the finite mixture model, where the component membership is implicit in the mixing weights $\pi_{k}$ and assumed to be binary (fully in one component). In the partial membership context, $\tau_{ik}$ as an exponent softens this assumption, allowing $\bm{x}_{i}$ to “partially belong” to multiple components simultaneously, each to varying degrees expressed by $\tau_{ik}$.
The conditional density $\tau\mid x_{i}$ is a continuous distribution over the simplex, capturing the relative contributions of each component to explaining the observation $x_{i}$. For example, in a two-component mixture ($K=2$) with components $N(0,1)$ and $N(4,1)$, observing $x=2$ leads to a continuous posterior distribution over $\tau_{i}=(\tau_{i1},\tau_{i2})$. A mode of $(\tau_{i1},\tau_{i2})=(0.5,0.5)$ would indicate equal partial membership in both components. This is conceptually distinct from finite mixtures, where $\tau_{i}=(0,1)$ or $(1,0)$ reflects binary cluster membership with uncertainty. The PM model thus enables nuanced interpretations of partial membership. For instance in the proposed application, a football player with $\tau=(0.5,0.5)$ can be understood as splitting their playing style equally between two roles, such as striker and midfielder. This interpretation reflects the simultaneous contribution of both clusters, rather than uncertainty about membership.
@@
In Figure 1, the red plate indicates nodes repeated for each observation $i=1,\dots,N$. Within it, $\bm{\tau}_{i}$ denotes the membership vector for observation $i$, drawn from a Dirichlet distribution with hyperparameter $\bm{\delta}$. The deterministic node $\bm{\mu}_{i}$ combines $\bm{\tau}_{i}$ and the cluster-specific Poisson parameters $\bm{\lambda}$ (enclosed in the blue plate), thereby defining the rates that generate the observed counts $\bm{X}_{i}$. Each element $\lambda_{k,j}$ in $\bm{\lambda}$ has a $\mbox{Gamma}(\alpha,\beta)$ prior, and arrows show how these parameters interact to govern the Poisson likelihood of the data. The nodes $(a,b)$ represent hyperparameters defining a $\mbox{Uniform}(0,10)$ prior for $\bm{\delta}$, while $(\alpha,\beta)$ serve as hyperparameters for the Gamma priors on $\bm{\lambda}$. Hence, all observations share a global prior structure but each observation’s partial membership vector $\bm{\tau}_{i}$ ultimately shapes $\mu_{i,j}$ and thus the distribution of $\bm{X}_{i}$.
While the Poisson independence assumption for the form of the distribution for each cluster may appear restrictive, it keeps the model tractable and interpretable. In particular, because partial membership already introduces a continuous mixing over clusters, adding a correlated structure for the Poisson variables can complicate inference and obscure each cluster’s “mean profile.” As a result, we focus on independent Poisson likelihoods, acknowledging that more flexible covariance structures could be explored in future work.
@@ 2.3 Mixed Membership Model
A method for fitting mixed membership (MM) models to count data is outlined in White and Murphy (2016). In broad terms, an MM model stipulates that each observation (here, each player) has a membership vector $\bm{\tau}_{i}$ drawn from a Dirichlet distribution. However, crucially, each attribute of observation $i$ (for instance, each type of pass or shot count) is then drawn from one of the $K$ latent clusters based on $\bm{\tau}_{i}$.
Formally, once $\bm{\tau}_{i}$ is obtained, a latent cluster label $\bm{Z}_{ij}$ is sampled for each attribute $j$, typically via a multinomial or categorical distribution with parameters $\bm{\tau}_{i}$. Next, conditional on $\bm{Z}_{ij}$, the attribute $x_{ij}$ is drawn from the corresponding cluster-specific distribution $P(\cdot\mid\lambda_{Z_{ij},j})$. This means that two attributes of the same player $i$ can come from two different clusters, reflecting a “mixed” or pattern at the attribute level.
In other words, MM models operate under the assumption that each data feature (such as pixels in an image analysis scenario) of a given item (e.g., an image) is conditionally independent, given the membership vector $\bm{\tau}_{i}$.
The likelihood for an observation is given by:
$x_{ij}\mid\bm{\tau}_{i}\sim\sum_{k}\tau_{ik}P({x}_{ij}\mid\lambda_{kj})\quad\mbox{for}\quad j=1,2,\ldots,J.$
This formulation relies on the assumption of “factorization over attributes,” meaning that the likelihood of an observation is expressed as the product of likelihoods for its individual attributes, given $\bm{\tau}_{i}$:
$P(x_{ij}\mid\bm{\tau}_{i})=\prod_{j=1}^{J}\sum_{k=1}^{K}\tau_{ik}P({x}_{ij}\mid\lambda_{kj}).$
While this simplifies the modeling process and is effective in capturing attribute-level patterns, it assumes conditional independence between attributes within an observation. This assumption holds if the attributes within a single observation are conditionally independent given $\bm{\tau}_{i}$. For example, in applications like image analysis, where pixels can be viewed as exchangeable sub-objects, this assumption is reasonable. In contrast, Partial Membership (PM) models relax the factorization assumption, allowing for joint modeling of attributes. This enables the PM model to account for attribute dependencies that may not align strictly with cluster-specific distributions.
@@
The key difference in the MM approach is its per-attribute labeling via $Z_{ij}$: a single observation can thus exhibit multiple latent roles across its various attributes, rather than having one role (as in a standard mixture) or a single partial membership vector that weights the entire attribute vector (as in PM).
@@ 2.5 Identifiability
Parameters in a model are not identified if the same likelihood function is obtained for more than one choice of the model parameters (Teicher, 1963; Gelman et al, 2013). The issue of identifiability in finite mixture distributions demands careful consideration, particularly in a Bayesian context. Frühwirth-Schnatter (2006) describes three sources of non-identifiability: (i) invariance to relabeling of components, (ii) non-identifiability stemming from potential overfitting, and (iii) a generic non-identifiability in certain classes of mixture distributions. Any of these can lead to poor convergence of MCMC methods.
In partial membership models, multiple sets of parameter values can again produce the same likelihood, creating further potential for non-identifiability. To address this, we leverage archetypal units, observations with near-complete membership in a single cluster. By serving as “extreme” examples, these units effectively anchor the corners of the membership simplex, thus stabilizing cluster definitions and enhancing interpretability. Although our terminology refers to “archetypal units,” we do not perform a separate archetypal analysis as in Cutler and Breiman (1994), nor do we create “prototype” clusters at the model level (as in Hou-Liu and Browne (2022a)). Rather, we identify archetypal units post hoc by selecting observations whose maximum membership weight is close to one. Alongside identifying archetypal units, we must also handle invariance to relabeling (label switching). Among the various solutions (e.g., Celeux et al, 2000; Murphy et al, 2020), we address label switching post hoc via a probabilistic relabeling algorithm (Sperrin et al, 2010), implemented in the R package label.switching. This approach permutes cluster labels in each MCMC draw so that the final posterior summaries remain label-consistent across iterations.
In our application, archetypal units serve a dual function. First, they enhance interpretability by identifying archetypes that mark the boundaries of each cluster, ensuring that the resulting partitions remain meaningful in the context of football performance. Second, they help mitigate the non-identifiability that is common in PM models by introducing tangible constraints on the cluster definitions. Thus, identifying these archetypal units makes the PM model more practically useful, particularly in domains where clear-cut examples are important for understanding cluster structure. By anchoring the model with these near-pure members (e.g., particular football players), the model can still capture mixed attributes for other players while avoiding overly diffuse or indistinct clusters.
@@
The selection of a uniform distribution with parameters $(0,10)$ as the hyperprior for the $\bm{\delta}$ parameter in the Dirichlet distribution governing the mixture weights, i.e., $\tau_{i}\sim\mathrm{Dirichlet}(\bm{\delta})$, proved instrumental in enabling the identification of archetypal players while preserving interpretability in the application in Section 3. For completeness, we also experimented with a sparser mixture-weight matrix by using a smaller scalar parameter in the Dirichlet (Wang and Blei, 2009), but found that this increased sparsity led to a higher optimal number of clusters based on our information criterion. Consequently, the uniform distribution emerged as our preferred option, effectively promoting the detection of archetypal units while keeping the cluster count manageable, thus preserving interpretability. Striking a middle ground, $\mbox{Uniform}(0,10)$ allowed some observations to achieve nearly complete membership in one cluster, effectively anchoring the membership simplex and thereby improving interpretability.